ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

11. Mathematical Physics

1991 MS Classification: Pflaum, Markus, Mathematisches Institut, Humboldt-Universität Remarks on the quantization of stratified symplectic manifolds We introduce a notion of deformation quantization for stratified symplectic spaces which extends the well-known deformation quantization on manifolds to singular spaces. For certain reduced spaces of circle actions on symplectic manifolds we show that a deformation quantization in the above sense exists indeed.